# Human2Robot: Learning Robot Actions from Paired Human-Robot Videos

### v1

@@ Abstract
Distilling knowledge from human demonstrations is a promising way for robots to learn and act. Existing work often overlooks the differences between humans and robots, producing unsatisfactory results. In this paper, we study how perfectly aligned human-robot pairs benefit robot learning. Capitalizing on VR-based teleportation, we introduce H&R, a third-person dataset with 2,600 episodes, each of which captures the fine-grained correspondence between human hands and robot gripper. Inspired by the recent success of diffusion models, we introduce Human2Robot, an end-to-end diffusion framework that formulates learning from human demonstrates as a generative task. Human2Robot fully explores temporal dynamics in human videos to generate robot videos and predict actions at the same time. Through comprehensive evaluations of 8 seen, changed and unseen tasks in real-world settings, we demonstrate that Human2Robot can not only generate high-quality robot videos but also excel in seen tasks and generalize to unseen objects, backgrounds and even new tasks effortlessly.
@@ 1 Introduction
The ability to observe others, whether people or animals, and learn skills to solve new tasks is a key reason humans can handle a wide range of complex challenges. To develop robots capable of assisting with various everyday problems, they must also acquire this skill, particularly the ability to learn directly from humans. This motivates a plethora of work studying how to effectively learn from human demonstrations [5, 58, 39, 59, 4, 55]. However, due to the discrepancies between humans and robots, existing work [59, 39, 70] struggles to transfer learned representations from humans to robots, or simply enforcing consistent outputs between humans and robots [4] without considering the motion trajectories.
@@
In this paper, we argue that fine-grained alignment between humans and robots is of great significance. This is because coarse-grained data provides limited clues for key points and cannot capture detailed information from humans, such as different movement trajectories or grasping positions. However, annotating the detailed correspondence between humans and robots, while feasible [29, 35], is quite time-consuming and labor-intensive. Fortunately, virtual reality (VR)-based teleoperation [28, 15, 45, 13] offers a promising way for collecting synchronized actions of humans and robots thanks to its efficiency and affordability. This motivates us to build upon VR-based teleoperation to study how paired data can help robots learn better. In particular, we focus on obtaining perfectly aligned data between human hands and robot arms as a testbed for learning from human demonstrations. With this in mind, we first introduce a data collection pipeline for gathering paired videos of human hands and robotic arms, aligned frame by frame. Compared to previous work that solely collects demonstration data, our method does not require much additional setup, equipment, or time, making it exceptionally easy to implement. It allows for efficient and consistent data collection. Furthermore, this approach can theoretically be extended to collect data from any type of robotic arms and a wide variety of scenes. Finally, we introduce H&R, a third-person dataset characterized by perfectly aligned videos of human hands and robotic arms, which comprises six fundamental tasks and six long-horizon tasks, totalling 2,600 episodes. Our goal is to train a robot arm that is able to mimic the actions performed by human hands and then perform tasks correspondingly, conditioned on a human demonstration. This requires the robot to know how its arm is aligned with human hands, which is readily available in our dataset, but more importantly, has the ability to act correspondingly. Therefore, the robot needs to predict what actions to take in the future conditioned on human motions as well as historical clues. This is similar in spirit to video generation tasks, which are able to produce perceptually-convincing future frames (actions) based on on a variety of clues [14, 23, 24, 25, 38, 42]. As a result, we build upon video generation models to mimic human actions.
@@
In light of this, we introduce Human2Robot, an end-to-end generative framework that predicts robotic videos and actions to take simultaneously. In particular, Human2Robot is built upon a pretrained Stable Diffusion model, including a spatial UNet for feature extraction, two behavior extractors for motion and position encoding, a spatial-temporal UNet that leverages temporal clues for frame and action prediction. As a result, Human2Robot is not only able to generate high-quality robotic arm videos, which is crucial for developing reliable and explainable models, as it enables users to visually verify and understand the underlying logic of robotic action trajectories, but can also predict corresponding action trajectory required to perform the task. With carefully designed training strategies, Human2Robot not only performs well on routine tasks but also generalizes to different positions, unseen objects, backgrounds and new tasks without additional training. Furthermore, we also introduce a KNN-based approach, enabling Human2Robot to perform routine tasks even without human demonstrations. In summary, our main contributions include:
@@
  • We introduce a low-cost, low-requirement, high-efficiency, and highly scalable pipeline for collecting perfectly aligned data of humans and robots in each frame, and we present H&R, the first dataset featuring perfectly aligned videos of humans and robots.
@@
  • We introduce Human2Robot, a diffusion-based model capable of generating robotic videos from human videos while simultaneously predicting the corresponding robotic trajectories in an end-to-end manner.
  • We propose a KNN+Human2Robot method, which leverages KNN to perform tasks without explicitly requiring human hand videos as a condition.
  • Human2Robot can generate high-quality robotic arm videos, which are crucial for developing reliable and explainable models, as well as excelling in routine tasks, even with variations in position, environment, and objects. Notably, Human2Robot can complete new tasks without any additional training and is capable of executing all tasks in real time.
@@ 2 Related Work
Teleoperation. Since current learning-based methods require large amounts of demonstration data, researchers have developed various forms of teleoperation approaches for data collection [34, 21, 54, 45, 57, 2, 3, 47, 19, 71, 17, 1]. However, these approaches may entail specific device requirements and may only be applicable to specific forms of robotic systems. Recently, VR-based methodologies [28, 15] have attracted considerable attention due to their cost-effectiveness, efficiency and versatility. However, these methods map human action information to robot operations and record the results, often overlooking the importance of the paired visual information. We leverage this perception to capture perfectly aligned videos of humans and robots, which are crucial for robotic imitation.
@@
Learning from Human Videos. Owing to the prohibitive expenses of collecting data on real robots, researchers have recently been attempting to leverage existing human-centric video datasets to enhance robot policy learning [35, 56, 10, 53, 70]. Researchers propose to learn representations from human videos to assist robots in task execution [64, 59, 37]. However, these approaches need substantial prior knowledge and struggle to transfer to robots [39, 5]. A recent approach [4] trains robots to make changes to the environment as humans do by simply enforcing consistent outputs between humans and robots. This could overlook crucial details provided by human hands including proper posture and feasible trajectory. Meanwhile, certain methods  [33, 29] have focused on aligning representations across videos, but they do not directly tackle the domain difference between human-centric videos and robot-centric videos. In contrast, our approach utilizes paired data and diffusion models to achieve precise alignment of human and robot videos.
@@
Diffusion Models for Video Generation. Due to the remarkable performance of diffusion models in image generation [48, 49, 41, 51, 6], many approaches have explored their potential for video generation [26, 16, 8, 63, 32, 66]. Moreover, conditionally controlled video generation [40, 30, 60, 27, 65] has broader and more practical application scenarios and has already become a hot topic in academic research. In recent work, several approaches [9, 62, 7] have utilized generative diffusion models to aid policy learning for robots. However, these methods primarily generate videos from a conventional human perspective, often overlooking the distributional differences between human and robotic actions. In contrast, we introduce a novel approach by applying video diffusion models to robotic scenarios, utilizing a generative model conditioned on human behaviors to guide robotic action outputs.
@@
Diffusion Models for Visual Understanding. The diffusion model not only performs impressively in generative domains but also shows promising applications in visual understanding tasks [69, 52, 36, 18]. Some notable approaches involve using diffusion models for object detection [11], image segmentation [67], and visual representation learning [69, 12]. The approach most similar to ours, GenRec [61] proposes a unified video diffusion model that jointly optimizes video generation and video classification, exhibiting the effectiveness of the diffusion model in video understanding. Our approach is designed for robotic tasks, constructing a human-to-robot video generative model that can simultaneously predict actions for subsequent robotic manipulation tasks.
@@ 3.1 Data Collection Pipeline
Our system setup is straightforward, consisting of a Meta Quest 3 VR headset, a 7-DoF xArm robotic arm, and two Intel Realsense D435 cameras capturing images of the human hand and robotic arm from the same viewpoint. The dual-camera setup ensures capturing the synchronized movements of both the human and robotic arms. With the integration of the hardware components, our system aims to achieve a critical objective:
Towards this goal, we build our H&R system focusing on developing software and establishing data collection protocols. We build the control system with OpenTeach [28]. When the user puts on the headset and performs several hand operations, the VR headset sends the coordinates of the hand movements of the user to the robotic arm and achieves synchronized actions. Here, one of the biggest challenges lies in the inconsistency of the scale and the orientation between the two coordinate systems. We find a simple yet effective solution to this problem, and the detailed method is provided in the Section B.
@@ 3.2 Statistics of H&R Dataset
With our designed system and data collection pipeline, we propose our H&R dataset, the first dataset featuring paired human and robotic videos. Human events vary from six fundamental tasks and six long-horizon tasks. The whole dataset includes 2,600 episodes, and each episode contains frames ranging from 300 to 600. Details can be found in the Section C.1.
To the best of our knowledge, the H&R dataset is the first video dataset that ensures perfectly aligned video between the hand and the robotic arm.
@@ 4 Human2Robot
Our goal is to develop a model capable of learning from human demonstrations to predict future robot observations and corresponding actions. Formally, we aim to establish a universal policy $\pi$ that predicts future robot observations $\hat{o}_{r}^{t:t+n}$ and action trajectories $a_{t:t+n}$, leveraging inputs from a sequence of hand demonstration videos $o_{h}^{t:t+n}$, environmental observations $o_{r}^{t}$, and the current proprioception $s_{t}$:
$a_{t:t+n},\ \hat{o}_{r}^{t:t+n}=\pi(o_{h}^{t:t+n},o_{r}^{t},s_{t}),$ (1)
@@ 4.1 Architecture
Inspired by the recent success of video diffusion models that are able to predict high-fidelity future frames [20], Human2Robot formulates the behavior learning process from human demonstrations as a generative task—generating robot videos and their corresponding actions simultaneously, conditioned on human videos and historical context. Inspired by  [27], Human2Robot explores a Spatial UNet and a Spatial-Temporal UNet, working collaboratively to learn from humans. In particular, the Spatial UNet extracts features from the robot arm, which is further fed into the Spatial-Temporal UNet for temporal modeling. Human2Robot also contains two behavior extractors that estimate position and motion clues using human videos. To achieve action prediction, we use the intermediate features from the Spatial-Temporal UNet to predict the action trajectory via a Policy Head. Figure 3 gives an overview of our pipeline. Below, we introduce the architecture in detail. Spatial UNet. The Spatial UNet, consisting of 4 upsampling blocks and 4 downsampling blocks, extracts useful clues from the robot arm. In particular, each block, known as the Spatial Layer, leverages self-attention to learn features tailored for robot arms, as well as cross-attention with pre-extracted CLIP embedding to incorporate semantic clues. The weights of the Spatial-UNet are initialized from those of the Stable Diffusion [50] to ease the learning process. The features derived by Spatial-UNet provide a condition for predicting future frames as well as actions.
@@
While Spatial UNet introduces a similar number of parameters as the denoising UNet, it only extracts features once during the entire process, unlike diffusion-based video generation where each frame undergoes denoising multiple times. Therefore, it does not significantly increase computational overhead during inference. Behavior Extractor. Recall that we aim to learn from human demonstrations, and thus it is important to extract motion and position information from human videos. This is achieved by Behavior Extractor, which contains four convolution layers (4$\times$ kernels, with a stride of 2). To focus more on the motion trajectories, we employ two independent Behavior Extractors. Behavior Extractor 1 takes the image of the hand corresponding to the robot at the first time step, and Behavior Extractor 2 takes the future image or the entire video of the hand as inputs depending on the training stage, as will be introduced in Section 4.2.
@@
Spatial-Temporal UNet. The Spatial-Temporal UNet aims to predict future frames as well as actions to take simultaneously, requiring the model to learn temporal dynamics. Thus, each block of the Spatial-Temporal UNet (ST-UNet in short) has a Spatial Layer which is the same as the one in Spatial UNet (S-UNet in short), followed by an additional Temporal Layer. More specifically, the ST-UNet takes the features from the behavior extractors and noise latent as inputs. Each Spatial Layer in ST-UNet also concatenates the features of the corresponding layer of the S-UNet, exploring the robot arm as references for prediction. The Temporal Layer focuses on temporal modeling so as to produce high-fidelity future frames as well as accurate action trajectories. Policy Head. We aim to generate future frames as well as actions to take within a single unified framework. Drawing from the insights of [61], we propose to divide the UNet mapping function into two distinct components: $F=F_{tail}\cdot F_{head}$. We treat $F_{head}$ as a feature extractor for motion information. This is achieved by passing the intermediate features from UNet and proprioception data $s_{t}$ to a light-weighted Multilayer Perceptron to predict the actions of the robotic arm. Pretrained Encoder and Decoder. Human2Robot builds upon Stable Diffusion [50], which consists of an encoder and a decoder. The encoder turns images into latent embeddings for fast denoising, and the decoder maps the latents back to images. During training, both the encoder and decoder are kept frozen.

### v2

@@ Abstract
Distilling knowledge from human demonstrations is a promising way for robots to learn and act. Existing work often overlooks the differences between humans and robots, producing unsatisfactory results. In this paper, we study how perfectly aligned human-robot pairs benefit robot learning. Capitalizing on VR-based teleportation, we introduce H&R, a third-person dataset with 2,600 episodes, each of which captures the fine-grained correspondence between human hand and robot gripper. Inspired by the recent success of diffusion models, we introduce Human2Robot, an end-to-end diffusion framework that formulates learning from human demonstration as a generative task. Human2Robot fully explores temporal dynamics in human videos to generate robot videos and predict actions at the same time. Through comprehensive evaluations of 4 carefully selected tasks in real-world settings, we demonstrate that Human2Robot can not only generate high-quality robot videos but also excels in seen tasks and generalizing to different positions, unseen appearances, novel instances, and even new backgrounds and task types.
@@ 1 Introduction
The ability to observe others, whether people or animals, and learn skills to solve new tasks is a key reason humans can handle a wide range of complex challenges. To develop robots capable of assisting with various everyday problems, they must also acquire this skill, particularly the ability to learn directly from humans. This motivates a plethora of work studying how to effectively learn from human demonstrations [5, 56, 38, 60, 4, 53]. However, due to the discrepancies between humans and robots, existing work [60, 38, 71] struggles to transfer learned representations from humans to robots, or simply enforce consistent outputs between humans and robots [4] without considering motion trajectories.
In this paper, we advocate that fine-grained alignment between humans and robots is of great significance. Coarse-grained data provides limited clues for key points and cannot capture detailed information, such as different movement trajectories or grasping positions. While annotating detailed correspondences between humans and robots is feasible [29, 34], it is time-consuming and labor-intensive. Fortunately, virtual reality(VR)-based teleoperation [28, 15, 43, 13] offers a promising way to collect synchronized actions of humans and robots, making it efficient and affordable. This motivates us to build upon VR-based teleoperation for studying paired data that can improve robot learning. In particular, we focus on obtaining perfectly aligned data between humans and robots as a testbed for learning from human demonstrations.
@@
To this end, we introduce a novel data collection pipeline for gathering paired videos of human hands and robotic arms, aligned frame by frame. Unlike previous work that solely collects demonstration data, our method requires minimal additional setup, equipment, or time, making it exceptionally easy to implement and facilitating efficient and consistent data collection. We alse introduce the H&R dataset, a third-person dataset featuring perfectly aligned videos of human hands and robotic arms, which includes 8 fundamental tasks and 6 long-horizon tasks, totalling 2,600 episodes. Our goal is to train a model that mimics human actions and performs corresponding tasks based on human demonstrations. This requires the robot to understand how its end effector is aligned with human hands, which is readily available in our dataset, but more importantly, has the ability to act correspondingly. Therefore, the robot needs to predict what actions to take in the future conditioned on human motions as well as historical clues. This is similar in spirit to video generation tasks, that predict future frames (actions) based on various cues [14, 23, 24, 25, 37, 41]. As a result, we build upon video generation models to mimic human actions.
@@
In light of this, we introduce Human2Robot, an end-to-end generative framework that predicts robotic videos and actions to take simultaneously. In particular, Human2Robot is built upon a pretrained Stable Diffusion model, including a spatial UNet for feature extraction, two behavior extractors for motion and position encoding, a spatial-temporal UNet that leverages temporal clues for frame and action prediction. As a result, Human2Robot is not only able to generate high-quality robotic videos, which facilitating the expansion of robot-conditioned policies to human-conditioned ones, but can also predict corresponding action trajectory required to perform the task. With carefully designed training strategies, Human2Robot not only excels on seen tasks but also generalizes to different positions, unseen appearances, instances and even new backgrounds and task types without additional training. Furthermore, we introduce a KNN-based approach, enabling Human2Robot to perform seen tasks even without human demonstrations, providing a flexible and scalable solution to robotic learning from demonstrations.
In summary, our main contributions include:
@@
  • We introduce a novel low-cost, low-requirement, high-efficiency, and highly scalable pipeline for collecting perfectly aligned data of humans and robots at the frame level. Additionally, we present H&R, the first dataset featuring perfectly aligned videos of human hands and robotic arms across a variety of tasks, enabling high-fidelity learning from human demonstrations.
@@
  • We introduce Human2Robot, an end-to-end generative framework built upon a diffusion-based model, capable of generating robotic videos from human videos while simultaneously predicting the corresponding robotic action trajectories. This approach combines state-of-the-art video generation techniques with multi-task action prediction, facilitating accurate task execution.
  • We propose a KNN+Human2Robot method, which integrates KNN for task prediction, enabling to perform tasks even without human hand videos as input. This further enhances the scalability and flexibility of the system.
  • Human2Robot can generate high-quality robotic arm videos, which facilitates the expansion of robot-conditioned policy to human-conditioned policy, while also excelling in four carefully selected tasks, even with variations in positions, appearances, instances, backgrounds and different task types.
@@ 2 Related Work
Teleoperation. Researchers have developed various forms of teleoperation approaches for data collection [33, 21, 52, 43, 55, 2, 3, 45, 19, 72, 17, 1]. However, these approaches may entail specific device requirements. Recently, VR-based methodologies [28, 15] have attracted considerable attention due to their cost-effectiveness, efficiency and versatility. However, these methods often overlook the importance of the paired visual information. We leverage this perception to capture perfectly aligned videos of humans and robots, which are crucial for robotic imitation.
@@
Learning from Human Videos. Researchers have recently been attempting to leverage existing human-centric video datasets to enhance robot policy learning [34, 54, 10, 51, 71]. Researchers propose to learn representations from human videos to assist in task execution [65, 60, 36]. However, these approaches need substantial prior knowledge and struggle to transfer to robots [38, 5, 4]. Meanwhile, some human-video-conditioned methods  [7, 29, 69] have focused on aligning representations across human and robot videos, they are neither efficient nor capable of generalization. In contrast, our approach utilizes paired data and diffusion models to achieve strong generalization capabilities.
@@
Diffusion Models for Video Generation. Due to the remarkable performance of diffusion models in image generation [46, 47, 40, 49, 6], many approaches have explored their potential for video generation [26, 16, 8, 64, 32, 67, 57, 58, 59]. Moreover, conditionally controlled video generation [39, 30, 61, 27, 66] has broader and more practical application scenarios. In recent work, several approaches [9, 63, 7] have utilized generative diffusion models to aid policy learning for robots. However, these methods primarily generate videos from a conventional human perspective, often overlooking the distributional differences between human and robotic actions. In contrast, we introduce a novel approach by applying video diffusion models to robotic scenarios, utilizing a generative model conditioned on human behaviors to guide robotic action outputs.
@@
Diffusion Models for Visual Understanding. The diffusion model not only performs impressively in generative domains but also shows promising applications in visual understanding tasks [70, 50, 35, 18]. Some notable approaches involve using diffusion models for object detection [11], image segmentation [68], and visual representation learning [70, 12]. The approach most similar to ours, GenRec [62] proposes a unified video diffusion model that jointly optimizes video generation and video classification, exhibiting the effectiveness of the diffusion model in video understanding. Our approach is designed for robotic tasks, constructing a human-to-robot video generative model that can simultaneously predict actions for manipulation tasks.
@@ 3.1 Data Collection Pipeline
Our system setup is straightforward, consisting of a Meta Quest 3 VR headset, a 6-DoF xArm robotic arm with a xArm gripper, and two Intel Realsense D435 cameras capturing images of the human hand and robotic arm from the same viewpoint.
The dual-camera setup ensures capturing the synchronized movements of both the human and robotic arms. With the integration of the hardware components, our system aims to achieve a critical objective: Producing a paired dataset where each data point reflects a seamless synchronization of human hand action and robotic arm response, captured synchronously from identical viewpoints.
Towards this goal, we build our H&R system focusing on developing software and establishing data collection protocols. We build the control system with OpenTeach [28]. When the user puts on the headset and performs several hand operations, the VR headset sends the coordinates of the hand movements of the user to the robotic arm and achieves synchronized actions.
Achieving this functionality requires addressing issues such as coordinate alignment, precision, and jitter. The details of the implementation are provided in the Supplementary materials B.2.
@@ 3.2 Statistics of H&R Dataset
With our designed system and data collection pipeline, we propose our H&R dataset, the first dataset featuring paired human and robotic videos. Human events vary from 8 fundamental tasks and 6 long-horizon tasks. The whole dataset includes 2,600 episodes, and each episode contains frames ranging from 300 to 600. Details can be found in Supplementary materials C.
To the best of our knowledge, the H&R dataset is the first video dataset that ensures perfectly aligned video between human and robot.
@@ 4 Human2Robot
Our goal is to develop a model capable of learning from human demonstrations to predict future robot observations and corresponding actions. Formally, we aim to establish a universal policy $\pi$ that predicts future robot observations $o^{r}_{t:t+n}$ and action trajectories $a_{t:t+n}$, leveraging inputs from a sequence of hand demonstration videos $o^{h}_{t:t+n}$, environmental observations $o^{r}_{t}$, and the current proprioception $s_{t}$:
$a_{t:t+n},\ o^{r}_{t:t+n}=\pi(o^{h}_{t:t+n},o^{r}_{t},s_{t}),$ (1)
@@ 4.1 Architecture
Inspired by the recent success of video diffusion models that are able to predict high-fidelity future frames [20], Human2Robot formulates the behavior learning process from human demonstrations as a generative task—generating robot videos and their corresponding actions simultaneously, conditioned on human videos and historical context. Inspired by  [27], Human2Robot explores a Spatial UNet and a Spatial-Temporal UNet, working collaboratively to learn from humans. In particular, the Spatial UNet extracts features from the robot arm, which is further fed into the Spatial-Temporal UNet for temporal modeling. Human2Robot also contains two behavior extractors that estimate position and motion clues using human videos. To achieve action prediction, we use the intermediate features from the Spatial-Temporal UNet to predict the action trajectory via a Policy Head. Figure 3 gives an overview of our pipeline. Below, we introduce the architecture in detail.
Spatial UNet. The Spatial UNet, consisting of 4 upsampling blocks and 4 downsampling blocks, extracts useful clues from the robot arm. In particular, each block, known as the Spatial Layer, leverages self-attention to learn features tailored for robot arms, as well as cross-attention with pre-extracted CLIP embedding to incorporate semantic clues. The weights of the Spatial-UNet are initialized from those of the Stable Diffusion [48] to ease the learning process. The features derived by Spatial-UNet provide a condition for predicting future frames as well as actions.
@@
While Spatial UNet introduces a similar number of parameters as the denoising UNet, it only extracts features once during the entire process, unlike diffusion-based video generation where each frame undergoes denoising multiple times. Therefore, it does not significantly increase computational overhead during inference.
Behavior Extractor. Recall that we aim to learn from human demonstrations, and thus it is important to extract motion and position information from human videos. This is achieved by Behavior Extractor, which contains four convolution layers (4$\times$ kernels, with a stride of 2). To focus more on the motion trajectories, we employ two independent Behavior Extractors. Behavior Extractor 1 takes the image of the hand corresponding to the robot at the first time step, and Behavior Extractor 2 takes the future image or the entire video of the hand as inputs depending on the training stage, as will be introduced in Section 4.2.
@@
Spatial-Temporal UNet. The Spatial-Temporal UNet aims to predict future frames as well as actions to take simultaneously, requiring the model to learn temporal dynamics. Thus, each block of the Spatial-Temporal UNet (ST-UNet in short) has a Spatial Layer which is the same as the one in Spatial UNet (S-UNet in short), followed by an additional Temporal Layer. More specifically, the ST-UNet takes the features from the behavior extractors and noise latent as inputs. Each Spatial Layer in ST-UNet also concatenates the features of the corresponding layer of the S-UNet, exploring the robot arm as references for prediction. The Temporal Layer focuses on temporal modeling so as to produce high-fidelity future frames as well as accurate action trajectories.
Policy Head. We aim to generate future frames as well as actions to take within a single unified framework. Drawing from the insights of [62], we propose to divide the UNet mapping function into two distinct components: $F=F_{tail}\cdot F_{head}$. We treat $F_{head}$ as a feature extractor for motion information. This is achieved by passing the intermediate features from UNet and proprioception data $s_{t}$ to a light-weighted Multilayer Perceptron to predict the actions of the robotic arm.
Pretrained Encoder and Decoder. Human2Robot builds upon Stable Diffusion [48], which consists of an encoder and a decoder. The encoder turns images into latent embeddings for fast denoising, and the decoder maps the latents back to images. During training, both the encoder and decoder are kept frozen.